OpenStudy (anonymous):

Solve for the variable. -22 = x - 7 A. x = -15 B. x = -29 C. x = 15 D. x = 29

OpenStudy (anonymous):

-22+7=x -15=x so A is the corect option

OpenStudy (anonymous):

@madrockz Thank you!!